# Transition to Petschek Reconnection in Subrelativistic Pair Plasmas: Implications for Particle Acceleration

### v1

@@ 4.3 Exhaust Spectrum
In the large-domain limit only a vanishingly small fraction of particles interact with the X-point at all. The vast majority are simply heated upon interaction with the exhaust. The basic mechanism is pickup acceleration as described in J. F. Drake et al. (2009): when a particle with negligible energy encounters a magnetohydrodynamic flow, the $E$ and $B$ fields accelerate the particle up to speed with the flow. This acceleration process is responsible for the spectral shoulder in Figure 7. Once within the flow, these accelerated particles are re-magnetized and undergo mirror motion along the curved field lines, bouncing between opposite boundaries of the exhaust. This behavior contrasts with recent results by N. K. Walia et al. (2022) concerning ions in electron-ion exhausts: there, due to the mass ratio, unmagnetized ions may gain energy reflecting off of shock potentials on either side of the exhaust. Such potentials are absent in pair plasma.
The net result of this particle energization is a thermal bulk flow proceeding at the Alfvén speed, with the energy distribution taking Maxwellian form in the comoving frame (Figure 11). Since the vast majority of the total dissipation occurs in these exhausts and not in X-points/plasmoids, we thus see that while high-$\sigma$ reconnection routinely generates strong power laws in energy, low-$\sigma$ reconnection merely heats the inflows (though a small, energetic contribution within the plasmoids is present). This is consistent with recent work emphasizing the importance of injection at X-points (L. Sironi, 2022; S. R. Totorica et al., 2023; S. Gupta et al., 2025): while the particles which happen to be injected there do reach nonthermal energies, those which do not will simply be part of the thermal bulk. Since injected particles are a minuscule part of the total for single X-point reconnection, heating dominates over nonthermal particle acceleration.
Figure 11: Fit of the observed spectra in a slice of the exhaust far from the X-point in Run V, where $\gamma^{\prime}$ is the energy as measured in the center-of-momentum frame. The spectrum is fitted to a Maxwell-Jüttner distribution with $\theta\equiv T/mc^{2}=0.06$. The discrepancy is due to incomplete velocity isotropization along the magnetic field.
@@ 5 Discussion and Conclusions
We then apply these results to explain the difference in observed particle spectra. For $\sigma<1$ the strong particle acceleration and power law formation associated with relativistic reconnection is replaced with thermal heating of particles through the pickup acceleration. We show via simulations and justify analytically that the X-points in the low-$\sigma$ regime are just as effective at accelerating particles as in the high-$\sigma$ regime: the reason they ultimately matter little is that few particles have the chance to interact with them. A tabular summary of the basic logic and conclusions is given in Table 3.
@@
Firstly, whether a reconnecting system proceeds as a plasmoid chain or as a Petschek configuration depends on the tearing rate. Either tearing produces sufficiently few plasmoids ($\mathcal{P}<\mathcal{P}_{c}$) that they uneventfully advect away from the X-point, resulting in a single X-line geometry, or it produces sufficiently many ($\mathcal{P}>\mathcal{P}_{c}$) that they undergo repeated mergers and thereby form a well-developed plasmoid chain. Hence, to predict whether a given system will, in the macroscopic limit, tend towards plasmoid-dominated or exhaust-dominated reconnection, one need only analyze the dynamics near a single X-point and compute $\mathcal{P}$. We expect the conclusions reached here for the pair plasma to also apply in the electron-ion regime: while mass ratio effects complicate the tearing physics, the same basic macroscale dichotomy remains. This framework can thus explain when one would expect to see an exhaust jet as opposed to a plasmoid chain in the heliosphere or in other environments.
@@
Secondly, the transition between the plasmoid- and exhaust-dominated reconnection is key for predicting particle acceleration. Although debate remains regarding the relative importance of direct $E$-field acceleration as opposed to Fermi mechanisms (e.g., F. Guo et al., 2025), our results demonstrate the necessity of plasmoid chains for power law generation. By allowing for stochastic acceleration at multiple X-points and within merging and compressing plasmoids, particles in relativistic reconnection can rapidly form a power-law distribution, avenues that are highly suppressed in an exhaust geometry. This has important effect on understanding whether nonrelativistic electron-ion reconnection is capable of generating strong power laws like those seen in simulations of high-$\sigma$ pair plasmas (e.g., X. Li et al., 2021). Since the injection physics is insensitive to magnetization (Section 4.2), we suggest that the answer will depend on whether the secondary tearing rate is high enough to generate frequent plasmoid mergers. Whether this would be observed in simulations depends on whether the box sizes are large enough to capture such dynamics.
@@
Lastly, our findings provide support for Petschek’s reconnection model. While there has been debate over whether Petschek reconnection can be realized from first-principles simulations (i.e., without ad hoc anomalous resistivity), here we demonstrate a simple case in which clear, stable laminar Petschek outflows form. While previous work suggested similar conclusions (M. Swisdak et al., 2008; Y.-H. Liu et al., 2012; M. E. Innocenti et al., 2015), ours extends these results in two main ways. Because we use outflow boundary conditions, we can wait for our simulations to reach a steady state that is neither a transient nor is sensitive to the initial conditions. Also, by simulating much larger domains (thousands of $c/\omega_{p}$), we confirm that plasmoid activity becomes less, not more, impactful as the box size is increased, meaning that these results can be extended to macroscopic system sizes. This bolsters the argument that the anomalous resistivity invoked in MHD simulations and theories of reconnection can indeed result from a fully self-consistent kinetic treatment. Ultimately, our results provide a physical framework for explaining the slow-mode shocks observed in the heliosphere as consequences of Petschek reconnection (e.g., N. K. Walia et al., 2024).
@@
We leave to further research the extension of these results to electron-ion plasmas and the important question of the role played by three-dimensional effects. Ultimately, the question of whether 3D reconnection proceeds as a dynamic multiple X-point phase or as a predominantly laminar single X-point configuration remains, although flux-rope breakdown and associated turbulence complicate the simple 2D picture (e.g., Q. Zhang et al., 2021).
The authors thank H. Ji, Y. H. Liu, N. Walia, and S. Zenitani for useful discussions. We thank R. Ewart for help with visualization of field lines. We are further grateful for H. Hakobyan’s assistance with TRISTAN V2. A.R. acknowledges support from NASA (80HQTR21T0105), and A.S. acknowledges the support of grants from NSF (PHY-2206607) and the Simons Foundation (MP-SCMPS-00001470).
Appendix A Derivation of Non-Relativistic Exit Spectrum
We consider a non-relativistic electron released from rest subject to $B_{x}=by$ and uniform $E_{z}$. The equations of motion are thus
$m\ddot{y}$ $=q\frac{\dot{z}}{c}by,$ (A1)
$m\ddot{z}$ $=-qE_{z}-\frac{q}{c}\dot{y}by.$ (A2)
@@ Appendix A Derivation of Non-Relativistic Exit Spectrum
$\ddot{y}$ $=\varepsilon y\dot{z},$ (A3)
$\ddot{z}$ $=-1-\varepsilon y\dot{y},$ (A4)
where $\varepsilon\equiv\tau_{E}\Omega$. For $\varepsilon\ll 1$, we can take $z=-t^{2}/2$ and use this to solve the resulting Airy equation Equation (A3)
$\ddot{y}=\varepsilon yt.$ (A5)
@@
$y=y_{0}A\left(-\frac{\,\text{Ai}\!\left({\zeta}\right)}{\,\text{Ai}^{\prime}\!\left({0}\right)}+\frac{\,\text{Bi}\!\left({\zeta}\right)}{\,\text{Bi}^{\prime}\!\left({0}\right)}\right),$ (A6)
@@
$A=\left(-\frac{\,\text{Ai}\!\left({0}\right)}{\,\text{Ai}^{\prime}\!\left({0}\right)}+\frac{\,\text{Bi}\!\left({0}\right)}{\,\text{Bi}^{\prime}\!\left({0}\right)}\right)^{-1}.$ (A7)
@@
$\,\text{Bi}\!\left({\zeta}\right)\approx\frac{1}{\sqrt{\pi}\zeta^{1/4}}\exp{\left(\frac{2}{3}\zeta^{3/2}\right)}.$ (A8)
@@
$y_{0}=y_{f}(2\sqrt{\pi}\,\text{Bi}\!\left({0}\right))\zeta^{1/4}\exp\left({-\frac{2}{3}\zeta^{3/2}}\right).$ (A9)
@@
$\frac{dN}{dW_{f}}=\frac{dN}{dy_{0}}\frac{dy_{0}}{dW_{f}},$ (A10)
where $W_{f}=t_{f}^{2}/2$ is the dimensionless kinetic energy when $y=y_{f}$. Substituting Equation (A9) and assuming $dN/dy_{0}$ is constant (i.e., a spatially uniform inflow density) yields our energy spectrum:
$\frac{dN}{dW_{f}}\propto W^{-1/4}\exp{\left(-\frac{2}{3}\varepsilon^{1/2}(2W)^{3/4}\right)}.$ (A11)
@@
Appendix B Hall-like Fields in the Diffusion Region

### v2

@@ 4.3 Exhaust Spectrum
In the large-domain limit only a vanishingly small fraction of particles interact with the X-point at all. The vast majority are simply heated upon interaction with the exhaust. The basic mechanism is pickup acceleration as described in J. F. Drake et al. (2009): when a particle with negligible energy encounters a magnetohydrodynamic flow, the $E$ and $B$ fields accelerate the particle up to speed with the flow. The resultant $E\times B$ cycloid motion imparts a total mean kinetic energy of $m(cE/B)^{2}=mv_{A}^{2}$, equally partitioned between the gyromotion and the guiding-center motion. This acceleration process is responsible for the spectral shoulder in Figure 5, with the mean pickup energy indicated by the dotted line. Once within the flow, these accelerated particles are re-magnetized and undergo mirror motion along the curved field lines, bouncing between opposite boundaries of the exhaust.
The net result of this particle energization is a thermal bulk flow proceeding at the Alfvén speed, with the energy distribution taking Maxwellian form in the comoving frame with $\theta\equiv T/mc^{2}=0.06$. Since the vast majority of the total dissipation occurs in these exhausts and not in X-points/plasmoids, we thus see that while high-$\sigma$ reconnection routinely generates strong power laws in energy, low-$\sigma$ reconnection merely heats the inflows (though a small, energetic contribution within the plasmoids is present). This is consistent with recent work emphasizing the importance of injection at X-points (L. Sironi, 2022; S. R. Totorica et al., 2023; S. Gupta et al., 2025): while the particles which happen to be injected there do reach nonthermal energies, the overwhelming majority which do not will simply be part of the thermal bulk.
@@ 5 Discussion and Conclusions
We then apply these results to explain the difference in particle spectra. For $\sigma<1$ the strong particle acceleration and power law formation associated with relativistic reconnection is replaced with thermal heating of particles through pickup acceleration. We show via simulations and justify analytically that the X-points in the low-$\sigma$ regime are just as effective at accelerating particles as in the high-$\sigma$ regime: the reason this ultimately matters little is that few particles have the chance to interact with them. A tabular summary of the basic logic and conclusions is given in Table 2.
@@
Firstly, whether a reconnecting system proceeds as a plasmoid chain or as a Petschek configuration depends on the tearing rate. Either tearing produces sufficiently few plasmoids ($\mathcal{P}<\mathcal{P}_{c}$) that they uneventfully advect away from the X-point, resulting in a single X-line geometry, or it produces sufficiently many ($\mathcal{P}>\mathcal{P}_{c}$) that they undergo repeated mergers and thereby form a well-developed plasmoid chain. Hence, to predict whether a given system will, in the macroscopic limit, tend towards plasmoid-dominated or exhaust-dominated reconnection, one need only analyze the dynamics near a single X-point and compute $\mathcal{P}$. We expect the conclusions reached here for pair plasma to also apply in the electron-ion regime: while mass ratio effects complicate the tearing physics, the same basic macroscale dichotomy remains. This framework can thus explain when one would expect to see an exhaust jet as opposed to a plasmoid chain in the heliosphere or in other environments.
@@
Secondly, the transition between the plasmoid- and exhaust-dominated reconnection is key for predicting particle acceleration. Although debate remains regarding the relative importance of direct $E$-field acceleration as opposed to Fermi mechanisms (e.g., F. Guo et al., 2025; O. French et al., 2025), our results demonstrate the necessity of plasmoid chains for power law generation. By allowing for stochastic acceleration at multiple X-points and within merging and compressing plasmoids, particles in relativistic reconnection can rapidly form a power-law distribution, avenues that are highly suppressed in an exhaust geometry. This has important effect on understanding whether nonrelativistic electron-ion reconnection is capable of generating strong power laws like those seen in simulations of high-$\sigma$ pair plasmas (e.g., X. Li et al., 2021). Since the injection physics is insensitive to magnetization (Section 4.2), we suggest that the answer will depend on whether the secondary tearing rate is high enough to generate frequent plasmoid mergers. Whether this would be observed in simulations depends on whether the box sizes are large enough to capture such dynamics.
@@
Lastly, our findings provide support for Petschek’s reconnection model. While there has been debate over whether Petschek reconnection can be realized from first-principles simulations (i.e., without ad hoc anomalous resistivity), here we demonstrate a simple such case in which clear, stable laminar Petschek outflows form. While previous work suggested similar conclusions (M. Swisdak et al., 2008; Y.-H. Liu et al., 2012; M. E. Innocenti et al., 2015), ours extends these results in two main ways. Because we use outflow boundary conditions, we can wait for our simulations to reach a steady state, neither a transient nor sensitive to the initial conditions. Also, by simulating much larger domains (thousands of $c/\omega_{p}$), we confirm that plasmoid activity becomes less, not more, impactful as the box size is increased, meaning that these results can be extended to macroscopic system sizes. This bolsters the argument that the anomalous resistivity invoked in MHD simulations and theories of reconnection can indeed result from a fully self-consistent kinetic treatment. Ultimately, our results provide a physical framework for explaining the slow-mode shocks observed in the heliosphere as consequences of Petschek reconnection (e.g., N. K. Walia et al., 2024).
@@
We leave to further research the extension of these results to electron-ion plasmas and the important question of the role played by three-dimensional effects. Ultimately, the question of whether 3D reconnection proceeds as a dynamic multiple X-point phase or as a predominantly laminar single X-point configuration remains, although flux-rope breakdown (e.g., Q. Zhang et al., 2021) and the drift-kink instability (e.g., F. Bacchini et al., 2025) complicate the simple 2D picture.
The authors thank H. Ji, Y. H. Liu, N. Walia, and S. Zenitani for useful discussions. We thank R. Ewart for help with visualization of field lines. We are further grateful for H. Hakobyan’s assistance with TRISTAN V2. We thank the anonymous reviewer for helpful suggestions which clarified the manuscript. A.R. acknowledges support from NASA (80HQTR21T0105), and A.S. acknowledges the support of grants from NSF (PHY-2206607) and the Simons Foundation (MP-SCMPS-00001470). This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
Appendix A Outflow Boundary Conditions
The outflow boundary conditions smoothly transition the electric and magnetic field components to target values near the top and bottom of the domain according to
$A$ $=e^{-\lambda}A_{\text{boundary}}+(1-e^{-\lambda})A_{\text{target}},$ (A1)
$\lambda$ $\equiv\kappa\left|(y-y_{\text{boundary}})/\Delta_{\text{abs}}\right|^{3},$ (A2)
where $A$ is a component of either the electric or magnetic field, $\Delta_{\text{abs}}$ is the absorption length, and $\kappa$ is a strength parameter. For a distance of $\Delta_{\text{abs}}$ cells from the domain edges in the $y$ direction the fields are set to the target value, and the transition occurs over another $\Delta_{\text{abs}}$ from that fixed region. For the electric field the target is $(0,0,0.1B_{0})$, and for the magnetic field the target is the initial value thereof. In our simulations, $\kappa=10$ and $\Delta_{\text{abs}}=10$.
Appendix B Derivation of Nonrelativistic Exit Spectrum
We consider a nonrelativistic electron released from rest subject to $B_{x}=by$ and uniform $E_{z}$. The equations of motion are thus
$m\ddot{y}$ $=q\frac{\dot{z}}{c}by,$ (B1)
$m\ddot{z}$ $=-qE_{z}-\frac{q}{c}\dot{y}by.$ (B2)
@@ Appendix B Derivation of Nonrelativistic Exit Spectrum
$\ddot{y}$ $=\varepsilon y\dot{z},$ (B3)
$\ddot{z}$ $=-1-\varepsilon y\dot{y},$ (B4)
where $\varepsilon\equiv\tau_{E}\Omega$. For $\varepsilon\ll 1$, we can take $z=-t^{2}/2$ and use this to solve the resulting Airy equation Equation (B3)
$\ddot{y}=\varepsilon yt.$ (B5)
@@
$y=y_{0}A\left(-\frac{\,\text{Ai}\!\left({\zeta}\right)}{\,\text{Ai}^{\prime}\!\left({0}\right)}+\frac{\,\text{Bi}\!\left({\zeta}\right)}{\,\text{Bi}^{\prime}\!\left({0}\right)}\right),$ (B6)
@@
$A=\left(-\frac{\,\text{Ai}\!\left({0}\right)}{\,\text{Ai}^{\prime}\!\left({0}\right)}+\frac{\,\text{Bi}\!\left({0}\right)}{\,\text{Bi}^{\prime}\!\left({0}\right)}\right)^{-1}.$ (B7)
@@
$\,\text{Bi}\!\left({\zeta}\right)\approx\frac{1}{\sqrt{\pi}\zeta^{1/4}}\exp{\left(\frac{2}{3}\zeta^{3/2}\right)}.$ (B8)
@@
$y_{0}=y_{f}(2\sqrt{\pi}\,\text{Bi}\!\left({0}\right))\zeta^{1/4}\exp\left({-\frac{2}{3}\zeta^{3/2}}\right).$ (B9)
@@
$\frac{dN}{dW_{f}}=\frac{dN}{dy_{0}}\frac{dy_{0}}{dW_{f}},$ (B10)
where $W_{f}=t_{f}^{2}/2$ is the dimensionless kinetic energy when $y=y_{f}$. Substituting Equation (B9) and assuming $dN/dy_{0}$ is constant (i.e., a spatially uniform inflow density) yields our energy spectrum:
$\frac{dN}{dW_{f}}\propto W^{-1/4}\exp{\left(-\frac{2}{3}\varepsilon^{1/2}(2W)^{3/4}\right)}.$ (B11)
@@
Appendix C Hall-like Fields in the Diffusion Region